OpenStudy (anonymous):

What is the degree of the following polynomial? Enter the numeric answer only. -9x^4+3x^2-7

OpenStudy (ledah):

here is an example on how to find the degree of a polynomial. Definition: The degree is the term with the greatest exponent Recall that for y2, y is the base and 2 is the exponent Example #1: 4x2 + 6x + 5 This polynomial has three terms. The first one is 4x2, the second is 6x, and the third is 5 The exponent of the first term is 2 The exponent of the second term is 1 because 6x = 6x1 The exponent of the third term is 0 because 5 = 5x0 What? 5x0 = 5? Well, anything with an exponent of 0 is always equal to 1 Thus, 5x0 = 5 × x0 = 5 × 1 = 5 Since the highest exponent is 2, the degree of 4x2 + 6x + 5 is 2 Hope you find this useful.

OpenStudy (cwrw238):

degree is value of highest exponent

OpenStudy (ledah):

so in this case. your highest exponent is 4.

OpenStudy (anonymous):

so its 4?

OpenStudy (ledah):

okay now try this one

OpenStudy (anonymous):

ok

OpenStudy (ledah):

2y6 + 1y5 + -3y4 + 7y3 + 9y2 + y + 6

OpenStudy (ledah):

what would be the highest exponent in that polynomial?

OpenStudy (anonymous):

6

OpenStudy (ledah):

2y^6 + 1y^5 + -3y^4 + 7y^3 + 9y^2 + y + 6 *revise*

OpenStudy (ledah):

very good. congrats you got it down